Solve for X
4/x+3 - 2/x-3 = 2

Answer by stanbon(75887)

Solve for X
4/x+3 - 2/x-3 = 2
---
Mulriply through by (x+3)(x-3) ??
4(x-3) - 2(x+3) = 2x^2-18
---
4x - 12 - 2x - 6 = 2x^2 - 18
2x^2 - 2x = 0
x^2 - x = 0
x(x-1) = 0
x = 0 or x = 1
======================
Cheers,
Stan H.